A cannon of mass 5m (including a shell of mass m) is at rest on a smooth horizontal ground ,fires the shell with its barrel at an angle theta with the horizontal at a velocity u relative to cannon.Find the horizontal distance of the pt where shell strikes the ground from the initial position of the canon.

Grade:11

1 Answers

Arun
5 years ago
since initial momentum is 0 final must also be,
relative=u
hence absolute=u + v...where v=recoil velocity,
 4m.v=m(u+v)
 v=u/3
hence initial velocity=u+u/3=4u/3
and angle of projection=A
hence Range=  16 u^2 sin2A
                        ____________    ANS
                             9g
 
Regards
Arun (a)skIITians forum expert